chris planted 1/6 of his watermelon seeds and 20 cucumber seeds. if he planted 26 seeds toal how many waermelon seeds does he have left?
Answer by checkley79(3341)

X/6=26-20
X/6=6
X=36 WATERMELON SEEDS LEFT.